1. [10pts] (a) Solve for 2 > 0 the DE: 20%Y + 102’y +59xy + 45y = 0. Answer: Solution: (b) Solve the DE in Part (a) for r
Answer #1

( 1 ) Yedico the (Oy x ²4 - 104 +59xy+ 45y = 0 linear romogeneous differential equation on into line arr differential equ7 (03+70°+519 +45)y=0 >> { DM9 +1) + 69 (0+1) +45(0+1)y = - }(0+1)(0760+45)}y = 0 Let y some be the trial solution of than th( 6 ) can The solution for a co be obtained from that of 270 by replacing n by L2 overyeruere. Thus, the solution of the dif